OpenStudy (anonymous):

Find a vector that the same direction as −2i + 4j+ 2k but that has length 6.

OpenStudy (anonymous):

desperately waiting for ur reply...so fed up with this problem right now !!! help

OpenStudy (anonymous):

hi guys

OpenStudy (anonymous):

do share ur views on this plss

OpenStudy (anonymous):

i am new here, on this site, looks fun

OpenStudy (anonymous):

surely it is !!! nothing lesser than a boon to the students !!! getting live help for free is just awesome

OpenStudy (anonymous):

feels like v r a community sharing its knowledge

OpenStudy (anonymous):

I am graduating highschool this spring

OpenStudy (anonymous):

congrats bro !!!

OpenStudy (anonymous):

the only bad thing is that me and my boyfriend are going to different universities :(

OpenStudy (anonymous):

which uni r u in ?

OpenStudy (anonymous):

Carnegie Mellon University

OpenStudy (anonymous):

what degree wer u in ?

OpenStudy (anonymous):

Computer Science, I know this sounds strange because I am a girl, but I LOVE computer science and programming

OpenStudy (anonymous):

thats good.....computers r not just monopoly of boyzzz

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

may i ask ur name ?

OpenStudy (anonymous):

ankur, you still need an answer?

OpenStudy (anonymous):

You just need to find the unit vector of u and multiply it by 6

OpenStudy (anonymous):

ah got it !!! cheers !!!

OpenStudy (anonymous):

Cool

OpenStudy (nikvist):

\[|-2i+4j+2k|=\sqrt{(-2)^2+4^2+2^2}=2\sqrt{6}\] \[\vec{a}=\alpha(-2i+4j+2k)\quad,\quad\alpha=\frac{6}{2\sqrt{6}}=\frac{\sqrt{6}}{2}\] \[vec{a}=\frac{\sqrt{6}}{2}(-2i+4j+2k)=-\sqrt{6}i+2\sqrt{6}j+\sqrt{6}k\]

OpenStudy (anonymous):

hey one more short question

OpenStudy (anonymous):

?

OpenStudy (anonymous):

unit vector that is perpendicular to the xz-plane

OpenStudy (anonymous):

(1,0,0) lies in x, (0,0,1) lies in z, so (0,1,0) will lie in y which is perpendicular to xz.

OpenStudy (anonymous):

@nikwist : could you please tell how do u type square roots like this on screen?

OpenStudy (anonymous):

and it's a unit vector.

OpenStudy (anonymous):

u r so intelligent !!! no wonder you have so many fans !!! thanx a lot again

OpenStudy (anonymous):

np...fan me if you haven't :)

OpenStudy (anonymous):

i have...do u have any idea how nikwist wrote his answers like that above?

OpenStudy (anonymous):

Yes, go into equation editor below, and there's a set of boxes to the right you can use. You can also type, sqrt{x} directly into the equation line to get\[\sqrt{x}\]

OpenStudy (anonymous):

cheers !!! u r really a lifesaver

OpenStudy (anonymous):

np :D